Find the vertices, eccentricity, foci and asymptotes of the hyperbola (x2/8) − (y2 /4) = 1. Also trace it. Under what conditions on λ the line x+λy = 2 will be tangent to this hyperbola? Explain geometrically.


1
Expert's answer
2021-06-12T05:23:24-0400

We compare this equation with "\dfrac{x^2}{a^2}-\dfrac{y^2}{b^2}=1."


Eccentricity is "e=\sqrt{1+\dfrac{b^2}{a^2}}=\sqrt{\dfrac{3}{2}}"


The center is "C=(0,0)"


The vertices are "V'=(−a,0)=(−2\sqrt{2},0)" and "V=(a,0)=(2\sqrt{2},0)"


To find the foci, we need the distance from the center to the foci "c^2=a^2+b^2=12, \,\,c=\pm 2\sqrt{3}"


The foci are "F'=(-c,0)=(2\sqrt{3},0)" and "F=(c,0)=(-2\sqrt{3},0)"


The asymptotes are "\dfrac{x^2}{8}-\dfrac{y^2}{4}=0,\,\, y=\pm \dfrac{1}{\sqrt{2}}x"





We compare equation of tangent to hyperbola "x_0\cdot\dfrac{x}{a^2}-y_0\cdot\dfrac{y}{b^2}=1" with "x+\lambda y=2"


We have "x-\dfrac{a^2}{b^2} \dfrac{y_0}{x_0}y=\dfrac{a^2}{x_0}, \,\, x-2\dfrac{y_0}{x_0}y=\dfrac{8}{x_0}" and "x_0=4", so "y_0=\pm2" and "\lambda=-\dfrac{y_0}{2}=\mp 1"